Be sure to answer all parts. Enter your answer in scientific notation. Calculate the equilibrium pressure of CO2 due to the decomposition of magnesium carbonate (MgCO3) at 25°C.

?=atm
Answer #1

the decompositin of MgCO3 can be represented as

MgCO3--------> MgO+ CO2

From lierature, the free energy change for the reaction = 64.87 Kj/mole at 298.15K

deltaG= -RTlnK

where K= Equilibrium constant

lnK= -detaG/RT= -64.87*1000/8.314*298.15=-7803

K= 4.31*10-12

for the given reaction K= [MgO] [PCO2]/ [MgCO3]

the activity of MgO and MgCO3=1

PCo2= 4.31*10-12 atm